# DiffGAN: A Test Generation Approach for Differential Testing of Deep Neural Networks

### v1

@@ 4.2 Baseline Approach
Since our main focus is to generate triggering inputs that reveal behavioral disagreements between pairs of DNN models, we selected the baseline approaches against which to compare our approach according to the following criteria.
@@
Finally, we exclude baselines that limit the generation of triggering inputs based on the number of initial test inputs available. Specifically, some approaches cap the maximum number of triggering images to match the size of a seed set—a small subset selected from the initial test dataset. This constraint can introduce bias into comparative evaluations. When methods are assessed within a fixed testing budget, often measured in terms of execution time, such limitations can lead to unfair performance comparisons. If one method is restricted by the number of initial inputs while others are not, it does not provide a level playing field for evaluation. Furthermore, in practical scenarios, the availability of initial test inputs might be limited. This scarcity makes it challenging to gauge a baseline method’s ability to generate additional triggering inputs beyond the initial set. From a practical standpoint, it is also difficult to accurately predict the maximum number of triggering inputs required to uncover all possible behavioral disagreements between the DNNs under test. Without knowing this number in advance, setting a cap based on initial seed inputs could hinder the thoroughness of the testing process.
@@
Based on the first criterion we identified DiffChaser [8], BET [11] Diverget [6], DRfuzz [10], and DFlare [9], as candidates. However, BET and Diverget were excluded due to their lack of open-source implementations. Additionally, DFlare was not selected because its approach to generating triggering inputs is constrained by the number of initial inputs in the testing dataset. The method applies multiple mutation rules to each initial test input to find a triggering input. Once a triggering input is found for a particular initial input, the method stops mutating that input and proceeds to the next one. This process limits the maximum number of possible triggering inputs to the number of initial inputs in the testing dataset, as it generates at most one triggering input per initial input. We also decided not to include DiffChaser in our experiments, as DRfuzz demonstrated superior performance in generating a greater number of trigger inputs [10]. Consequently, we only selected DRfuzz as a baseline to evaluate the effectiveness of our proposed approach.
@@
DRfuzz is an approach designed for generating test inputs that uncover diverse regression faults in updated versions of DNN models. In the context of DNNs, a regression fault occurs when an input, correctly predicted by a previous model version, leads to incorrect predictions in the updated version. DRfuzz initiates this process with seed inputs, upon which it applies various mutation rules to produce new inputs that trigger regression faults. To ensure the validity and relevance of these inputs, DRfuzz employs the discriminator component of a trained GAN to filter invalid inputs. A customized reward function in DRfuzz assigns priority scores to each mutation rule based on their effectiveness in generating diverse triggering inputs from the seed inputs. This selection strategy emphasizes mutation rules that are more likely to uncover behavioral disagreements between model versions. To align DRfuzz with our specific research objectives, we have customized the source code of DRfuzz accordingly. We refined the approach to keep not only the inputs that one model predicts correctly while being mispredicted by the other model, but also any inputs that show behavioral disagreements between the models under test.
4.3 Datasets and Models
@@
Pair Model Architecture Training Data Optimizer Parameters Accuracy
Pair 1 Model 1 4 Conv, 20 layers MNIST Org (60,000 images) Adam (lr=0.001) #Param: 330,730, Batch: 128, Epoch: 10 99.26%
Model 2 7 Conv, 19 layers MNIST Org (60,000 images) Adam (lr=0.001) #Param:: 327,242, Batch: 128, Epoch: 10 99.50%
@@
RMSprop (lr=0.001) #Param: 696,402, Batch: 128, Epoch: 10 99.09%
@@
RMSprop (lr=0.001) #Param: 696,402, Batch: 128, Epoch: 10 99.55%
Pair 3 Model 5 LeNet-5 MNIST Org (60,000 images) SGD (lr=0.01) #Param: 44,426, Batch: 32, Epoch: 10 97.04%
Model 6 LeNet-5 MNIST Org (60,000 images) Adam (lr=0.001) #Param: 44,426, Batch: 32, Epoch: 10 98.93%
Pair 4 Model 7 AlexNet MNIST Org (60,000 images) Adam (lr=0.00001) #Param: 24,733,898, Batch: 128, Epoch: 9 98.85%
Model 8 AlexNet MNIST Org (60,000 images) Adam (lr=0.0001) #Param: 24,733,898, Batch: 256, Epoch: 14 99.33%
@@
Pair Model Architecture Training Data Optimizer Parameters Accuracy
Pair 1 Model 1 ResNet18 Cifar-10 Org (60,000 images) SGD (lr = 1e-1) #Param: 11183562, Batch=128, Epoch=200 83.38%
Model 2 ResNet20 Cifar-10 Org (60,000 images) SGD (lr = 1e-1) #Param: 274442, Batch=128, Epoch=200 85.71%
@@
SGD (lr = 1e-1) #Param: 5984858, Batch=128 , Epoch=120 93.02%
@@
SGD (lr = 1e-1) #Param: 5984858, Batch=128, Epoch=120 93.27%
Pair 3 Model 5 Inception Cifar-10 Org (60,000 images) SGD (Nesterov=True) #Param: 5984858, Batch=128, Epoch=150 89.10%
Model 6 Inception Cifar-10 Org (60,000 images) Adam (lr = 1e-3) #Param: 5984858, Batch=128, Epoch=150 92.78%
@@
Figure 2: Examples of original MNIST images (considered valid).
@@
Additionally, we specifically focused on model pairs where the accuracy difference between models was minimal. It ensures that our evaluation captures subtle differences in model performance. Detecting such small variations is crucial for demonstrating that our framework can reliably identify differences even when models achieve nearly identical, high accuracy.
Figure 3: Examples of original Cifar-10 images.
@@ 5 Empirical Results
To answer this research question, we executed DiffGAN and DRfuzz considering the same execution time $T\in[1h,2h,3h]$, across four pairs of DNN models and two distinct datasets, resulting in eight unique combinations (Section 4.3). Because of the randomness in DiffGAN and DRfuzz, we re-executed each of them three times and reported the corresponding number of triggering inputs in Tables III and IV. This decision to limit the execution of the approaches to three iterations was initially due to the considerable execution time associated with the experiments. More specifically, our experiments take 96h (8 model pairs x 6h x 2 approaches) if we execute each approach once. As shown in the tables, DiffGAN consistently outperforms DRfuzz in terms of generating more triggering inputs across all models and datasets. More specifically, our results show that DiffGAN generates on average four times more triggering inputs than DRfuzz when considering the same testing budget. We also calculated the Coefficient of Variation (CV) [56], a measure of relative variability that compares the standard deviation to the mean, independent of the data’s scale. We computed the CV score for each model pair and testing budget across the three runs. A lower CV indicates higher consistency across different runs. For example, in our analysis of the MNIST dataset, DRfuzz showed a CV between 0.33% and 21.23%, while DiffGAN exhibited a CV between 1.33% and 12.20%, suggesting less variability in the number of generated triggering inputs across the different runs. These observations are also consistent for the Cifar-10 dataset. These CV values demonstrate the relative stability and reliability of DiffGAN in generating triggering images, with less variability compared to DRfuzz.
@@
Table III: RQ1 results: Number of generated triggering images for models trained on MNIST dataset using DiffGAN and DRfuzz
@@
Table IV: RQ1 results: Number of generated triggering images for models trained on Cifar-10 dataset using DiffGAN and DRfuzz
@@
To systematically compare DiffGAN and DRfuzz in terms of generating valid trigger inputs within a given time budget, we considered the inputs generated in RQ1 by both methods. More specifically, we considered all four model pairs using the MNIST dataset and run both approaches for two hours on these models. As mentioned in RQ1, we ran each method three times to minimize randomness and recorded all images generated by each approach during each run. After completing the runs, for each approach and model pair, we aggregated the images generated from all three runs into a comprehensive dataset. By combining the results from three separate runs for each configuration, we were able to mitigate the effects of randomness in the generation process, thereby increasing the reliability of our comparisons.
Table V: MNIST Improvement
@@
Table VI: Cifar-10 Improvement
@@
Selecting representative samples. Due to the large number of triggering inputs generated by DiffGAN and DRfuzz, we selected a sample of triggering inputs from each aggregated dataset. For each setting, we determined the sample size that can confidently represent the whole corresponding dataset with a 95% confidence level and a margin of error of 5% [^5] As computed from https://www.calculator.net/sample-size-calculator.html. As a result, in our experiments, the sample size varied from 361 to 381 across the four settings that we considered, as shown in Table VII. This approach helped choose a minimal subset that was large enough to provide results that accurately reflected the overall distribution of inputs generated by DiffGAN and DRfuzz [57].
Manual validity checking. From the non-redundant set, we randomly sampled the selected images for validation, according to the sample size determined the previous step . To ensure unbiased evaluation, we shuffled images generated by both DiffGAN and DRfuzz and presented them anonymously to three independent expert evaluators. These experts were tasked with assessing the validity of each image based on predefined criteria. Specifically, they labeled images as either valid or invalid. Invalid images were defined as those that displayed two numbers, no number, or other issues such as (1) Images that were a 50/50 blend of two digits and could not be clearly identified as a specific digit, (2) Incomplete or blended digits that could not be classified into one category by consensus, (3) Images containing random shapes or non-digit drawings, or (4) Images that were completely white or dark, with no discernible content. Examples of valid and invalid inputs generated by both approaches are depicted in Figure 4. To determine the validity of each image, we considered the agreement among the evaluators. If at least two out of three evaluators labeled an image as valid, we considered it valid. To quantify the level of agreement between the evaluators, we employ Cohen’s kappa [58]—a statistical measure that accounts for agreement occurring by chance. In our study, the value of Cohen’s kappa is 70%, indicating a high level of agreement among the evaluators and underscoring the reliability of the manual labeling process.
@@
Reporting the total number of valid inputs. We calculated the proportion of valid images for each method by dividing the number of valid images by the total number of sampled images. We then multiplied this validity ratio by the average number of images generated in the runs to estimate the expected validity ratio, the 95% confidence interval, and the number of valid images, as reported in Table VII.
Finally, we compared the expected number of valid inputs generated by DiffGAN and DRfuzz to assess which method was more effective at producing valid triggering inputs within the same time budget of two hours. The results presented in Table VII, clearly suggest that DiffGAN significantly outperforms DRfuzz in generating more valid triggering inputs. For instance, in the first setting, it is estimated that DiffGAN generated 8,292 valid triggering inputs, whereas DRfuzz produced only 1,999 such inputs. This corresponds to a validity ratio of 91.25% for DiffGAN compared to 82.46% for DRfuzz. Importantly, those results were consistent across all the settings we considered. To further analyze the statistical significance of our results we performed a statistical analysis using the Wilcoxon signed-rank test [60], with a significance level of $\alpha=0.05$, to assess whether DiffGAN significantly outperforms DRfuzz in generating valid triggering inputs. The Wilcoxon signed-rank test, a non-parametric method, is used to compare the medians of continuous variables for paired samples and is ideal for our case since it does not require any assumptions about the data distribution.
Instead of considering the number of valid inputs, we relied on the validity ratio due to differences in sample sizes. We evaluated the validity ratios from 4000 samples for DiffGAN and 4000 samples for DRfuzz, selected via bootstrapping across four different scenarios (1000 samples per scenario). The Wilcoxon signed-rank test, with w-statistic = 343465.5 and a p-value of 0.000001, confirmed that DiffGAN significantly outperformed DRfuzz at generating valid triggering inputs across all scenarios.
Table VII: Validity Comparison
@@
The results, as shown in Tables VIII and IX, consistently demonstrate that DiffGAN outperforms DRfuzz in generating more diverse triggering inputs, when considering both metrics, highlighting its effectiveness in exploring a broader range of input variations. Similarly to RQ1, we analysed the variability of our diversity results using the CV score. Our CV analysis shows that both DiffGAN and DRfuzz exhibit low variability across different runs, with DiffGAN demonstrating greater consistency. DRfuzz’s GD diversity has a CV range of 6.76% to 18.44%, while DiffGAN is more stable with a CV range of 2.24% to 14.23%. In terms of Shannon metrics, DRfuzz shows a CV range between 0.09% and 0.70%, whereas DiffGAN is even more consistent, with a CV range of 0.02% to 0.33%. These results underscore the superior stability of DiffGAN’s outputs compared to DRfuzz. Furthermore, the superior performance of DiffGAN in generating more diverse triggering inputs can be attributed to its large latent space, which allows for a greater variety of generated inputs, and its diversity-based fitness function specifically designed to maximize the diversity of generated inputs, a feature that is limited in DRfuzz. Indeed, DRfuzz focuses on triggering behavioral disagreements through limited mutation rules, which, while effective in some contexts, do not inherently prioritize or ensure diversity. As a result, the input sets generated by DRfuzz tend to be less diverse compared to DiffGAN.
Table VIII: Diversity scores of MNIST
@@
Table IX: Diversity scores of Cifar-10
@@
Figure 5: Accuracy of Random Forest in Selecting the Best Model
@@
We aim in this question to study whether we can better guide the training of an ML-based model selection mechanism that consists of (1) two classifiers (model 1 and model 2) responsible for image classification, and (2) a third model that dynamically selects which of the two classifiers to use based on the input characteristics, which is based on Random Forest (RF). It predicts, based on the characteristics of the incoming input, which of the two classifiers (model 1 or model 2) will produce a more accurate output, without the need to run both models in an online setting. When properly trained with triggering inputs, the RF model should learn when each classifier is more accurate, effectively serving as a decision mechanism for selecting the optimal classifier for a given input. We choose RF because: (1) it scales effectively to handle a large number of features, and (2) its robustness against overfitting [68, 69, 70]. We do not intend to provide a novel voting mechanism here but, instead, provide an example of an important application scenario for the triggering inputs we generate. It also serves as a basis to compare the value of the inputs generated by DiffGAN to enable such model selection and compare it with that of DRfuzz.
@@
To compare the accuracy of the RF-based model selection systems, we categorized the images into three groups based on whether one model predicted correctly or if both models failed. We excluded images where both models failed from further analysis due to their low occurrence and lack of relevance to the comparison (i.e., if both models fail, it does not matter which classification model is selected). We trained the RF classifiers using features extracted from a pre-trained VGG16 model, rather than raw pixel data, in an attempt to improve accuracy. The VGG16 model provides feature vectors that capture more abstract and informative patterns than pixel-based features, which is expected to improve the ability of the RF model to distinguish between correct and incorrect predictions. For each method and model pair, we selected 700 valid labeled images and used their VGG16 features to train eight RF-based selection models. To ensure a fair comparison between the RF models trained on different datasets in each scenario, we need a common test dataset. For each scenario, we selected 100 images from the DiffGAN-generated test set and 100 from the DRfuzz-generated test set, creating a combined test set of 200 images. This test set allowed us to evaluate the performance of the RF models using the same test set in an unbiased manner. Additionally, we applied the same feature extraction process as for the training sets.
We report the accuracy levels of each of the four selection models we trained in Figure 5. As clearly visible, the RF-based selection models trained using triggering inputs generated with DiffGAN are far more accurate than those trained with inputs from DRfuzz. For instance, in the first scenario involving two RF-based selection models built using classifiers Model 1 and Model 2 from Table I, the RF model trained with DRfuzz achieved an accuracy of 78%, whereas the one trained with DiffGAN reached a much higher accuracy of 87%. Overall, the accuracy levels for the RF models trained with DiffGAN range between 79% and 87% compared to a range of 70% to 78% for those with DRfuzz inputs. This highlights that DiffGAN produces more effective triggering inputs for training, allowing the RF model to better capture the conditions under which one classifier outperforms the other. As a result, RF models trained using DiffGAN inputs are more accurate in selecting the best-performing classifier, leading to a better model selection mechanism.

### v2

@@ 4.2 Baseline Approach
Selection Process. Since our main focus is to generate triggering inputs that reveal behavioral disagreements between pairs of DNN models, we selected the baseline approaches against which to compare our approach according to the following criteria.
@@
Finally, we exclude baselines that limit the generation of triggering inputs based on the number of initial test inputs available. Specifically, some approaches cap the maximum number of triggering images to match the size of a seed set—a small subset selected from the initial test dataset. This constraint can introduce bias into comparative evaluations. When methods are assessed within a fixed testing budget, often measured in terms of execution time, such limitations can lead to unfair performance comparisons. If one method is restricted by the number of initial inputs while others are not, it does not provide a level playing field for evaluation. Furthermore, in practical scenarios, the availability of initial test inputs might be limited [53, 8]. This scarcity makes it challenging to gauge a baseline method’s ability to generate additional triggering inputs beyond the initial set. From a practical standpoint, it is also difficult to accurately predict the maximum number of triggering inputs required to uncover all possible behavioral disagreements between the DNNs under test. Without knowing this number in advance, setting a cap based on initial seed inputs could hinder the thoroughness of the testing process.
@@
Based on the first criterion we identified DiffChaser [8], BET [11], Diverget [6], DRfuzz [10], and DFlare [9], as candidates. However, BET and Diverget were excluded due to their lack of open-source implementations. We also decided not to include DiffChaser in our experiments, as DRfuzz has demonstrated superior performance in generating a greater number of trigger inputs [10]. Consequently, we selected DRfuzz as a baseline to evaluate the effectiveness of our proposed approach. We should note that DFlare was excluded because its test generation process is inherently constrained by the number of initial test inputs. The method applies multiple mutation rules to each seed input but halts once a triggering input is found, preventing further mutation. As a result, DFlare can generate at most one triggering input per initial seed, significantly limiting the overall number of triggering inputs. To mitigate this limitation, we attempted to repeatedly feed the same seed inputs into DFlare to increase the number of generated inputs. However, due to DFlare’s low randomness in input generation, repeated executions produced highly redundant outputs. This redundancy is further amplified by the limited number of mutation operations and the exploitation of the same mutation operators that successfully increase behavioral disagreements between the models under test over the generations. As a result, when repeatedly fed with seed inputs, DFlare’s mutation process introduces limited variations, failing to produce a diverse set of triggering inputs. To quantify this limitation, we compared the diversity of the inputs generated by DFlare, DRFuzz, and our approach. Our results showed that DFlare’s generated inputs had significantly lower diversity scores than those produced by both DRFuzz and DiffGAN. Therefore, we excluded this baseline in our study.
@@
Selected Baseline. DRfuzz is an approach designed for generating test inputs that uncover diverse regression faults in updated versions of DNN models. In the context of DNNs, a regression fault occurs when an input, correctly predicted by a previous model version, leads to incorrect predictions in the updated version. DRfuzz initiates this process with seed inputs, upon which it applies various mutation rules to produce new inputs that trigger regression faults. To ensure the validity and relevance of these inputs, DRfuzz employs the discriminator component of a trained GAN to filter invalid inputs. A customized reward function in DRfuzz assigns priority scores to each mutation rule based on their effectiveness in generating diverse triggering inputs from the seed inputs. This selection strategy emphasizes mutation rules that are more likely to uncover behavioral disagreements between model versions. To align DRfuzz with our specific research objectives, we have customized the source code of DRfuzz accordingly. We refined the approach to keep not only the inputs that one model predicts correctly while being mispredicted by the other model, but also any inputs that show behavioral disagreements between the models under test.
@@
Pair Model Architecture Training Data Optimizer Parameters Accuracy # Mispredicted
Pair 1 Model 1 4 Conv, 20 layers MNIST Org (60,000 images) Adam (lr=0.001) #Param: 330,730, Batch: 128, Epoch: 10 99.26% 74
Model 2 7 Conv, 19 layers MNIST Org (60,000 images) Adam (lr=0.001) #Param:: 327,242, Batch: 128, Epoch: 10 99.50% 50
@@
RMSprop (lr=0.001) #Param: 696,402, Batch: 128, Epoch: 10 99.09% 91
@@
RMSprop (lr=0.001) #Param: 696,402, Batch: 128, Epoch: 10 99.55% 45
Pair 3 Model 5 LeNet-5 MNIST Org (60,000 images) SGD (lr=0.01) #Param: 44,426, Batch: 32, Epoch: 10 97.04% 296
Model 6 LeNet-5 MNIST Org (60,000 images) Adam (lr=0.001) #Param: 44,426, Batch: 32, Epoch: 10 98.93% 107
Pair 4 Model 7 AlexNet MNIST Org (60,000 images) Adam (lr=0.00001) #Param: 24,733,898, Batch: 128, Epoch: 9 98.85% 115
Model 8 AlexNet MNIST Org (60,000 images) Adam (lr=0.0001) #Param: 24,733,898, Batch: 256, Epoch: 14 99.33% 67
@@
Pair Model Architecture Training Data Optimizer Parameters Accuracy # Mispredicted
Pair 1 Model 1 ResNet18 Cifar-10 Org (60,000 images) SGD (lr = 1e-1) #Param: 11183562, Batch=128, Epoch=200 83.38% 1662
Model 2 ResNet20 Cifar-10 Org (60,000 images) SGD (lr = 1e-1) #Param: 274442, Batch=128, Epoch=200 85.71% 1429
@@
SGD (lr = 1e-1) #Param: 5984858, Batch=128 , Epoch=120 93.02% 698
@@
SGD (lr = 1e-1) #Param: 5984858, Batch=128, Epoch=120 93.27% 673
Pair 3 Model 5 Inception Cifar-10 Org (60,000 images) SGD (Nesterov=True) #Param: 5984858, Batch=128, Epoch=150 89.10% 1090
Model 6 Inception Cifar-10 Org (60,000 images) Adam (lr = 1e-3) #Param: 5984858, Batch=128, Epoch=150 92.78% 722
@@ 4.3 Datasets and Models
Figure 2: Original MNIST images (considered valid).
@@
Additionally, we specifically focused on model pairs where the accuracy difference between models was small. More specifically, the accuracy differences across all model pairs range from 0.24% to 3.68%. We further report in Tables I and II the number of mispredicted inputs for each model and highlight their small differences for each model pair. This ensures that our evaluation captures subtle differences in model performance. Detecting such small variations is crucial for demonstrating that our framework can reliably identify differences even when models achieve nearly identical, high accuracy.
Figure 3: Original Cifar-10 images (considered valid).
4.4 Choice of the Fitness Functions
To assess the contribution of each fitness function to the effectiveness of our approach in terms of generating diverse triggering inputs, we conducted an ablation study by isolating the impact of each fitness function used in our multi-objective search: divergence and diversity. We executed three configurations of our search algorithm:
  • Divergence Only: The search is guided solely by the divergence-based fitness function, which maximizes the disagreement between the output probability vectors of the DNN models under test.
  • Diversity Only: The search is guided solely by the diversity-based fitness function, which maximizes the diversity among the generated inputs.
  • Combined Objectives (Full DiffGAN): The full DiffGAN configuration using NSGA-II to optimize both fitness functions simultaneously.
Due to the computational expense of our experiments, we conducted this study on two model pairs with the Cifar-10 dataset. Seeing significant differences across configurations on such pairs would however be enough to justify the combined use of divergence and diversity. We studied both the total number of triggering inputs generated by each configuration and their diversity. To quantify diversity, we leveraged two metrics: Shannon Diversity [59, 60] and Geometric Diversity [1]. A detailed explanation of these metrics, along with our methodology for computing diversity scores, is provided in Section 5.3. The results of this ablation study are summarized in Table III.
Table III: Ablation Study Results for Two Model Pairs with Cifar-10
Models & Dataset Configuration GD Exponential Shannon ($e^{H}$) # Triggering Images
M1, M2 (Cifar-10) DiffGAN 2396.99 $e^{14.84}=2.79\times 10^{6}$ 80,700
Divergence Only 739.10 $e^{13.23}=5.57\times 10^{5}$ 109,146
Diversity Only 2656.00 $e^{14.94}=3.08\times 10^{6}$ 14,538
M3, M4 (Cifar-10) DiffGAN 2395.81 $e^{14.84}=2.79\times 10^{6}$ 62,596
Divergence Only 892.54 $e^{13.46}=7.01\times 10^{5}$ 90,811
Diversity Only 2652.34 $e^{14.94}=3.08\times 10^{6}$ 11,746
As shown in the table, we had consistent results for both model pairs. More specifically, our results show that using the divergence-based fitness function alone leads to a higher number of triggering inputs compared to the diversity-only configuration. However, these inputs often exhibit redundancy, which limits the exploration of diverse disagreement scenarios. On the other hand, the diversity-only configuration generates more diverse inputs but results in fewer behavioral disagreements between models. The full DiffGAN configuration, which optimizes both objectives simultaneously, provides a good trade-off between the number of generated triggering inputs and their diversity. Not only does it produce a substantial number of triggering inputs, but it also maintains high diversity among them. This confirms the complementary nature of the two fitness functions and underscores the importance of combining them to generate a diverse and informative set of triggering inputs within a limited testing budget.
@@ 5 Empirical Results
To answer this research question, we executed DiffGAN and DRfuzz considering the same execution time $T\in[1h,2h,3h]$, across four pairs of DNN models and two distinct datasets, resulting in eight unique combinations (Section 4.3). Because of the randomness in DiffGAN and DRfuzz, we re-executed each of them three times and reported the corresponding number of triggering inputs in Tables IV and V. This decision to limit the execution of the approaches to three iterations was initially due to the considerable execution time associated with the experiments. More specifically, our experiments take 96h (8 model pairs x 6h x 2 approaches) if we execute each approach once. As shown in the tables, DiffGAN consistently outperforms DRfuzz in terms of generating more triggering inputs across all models and datasets. More specifically, our results show that DiffGAN generates on average four times more triggering inputs than DRfuzz when considering the same testing budget. We also calculated the Coefficient of Variation (CV) [61], a measure of relative variability that compares the standard deviation to the mean, independent of the data’s scale. We computed the CV score for each model pair and testing budget across the three runs. A lower CV indicates higher consistency across different runs. For example, in our analysis of the MNIST dataset, DRfuzz showed a CV between 0.33% and 21.23%,while DiffGAN exhibited a CV between 1.33% and 12.20%, suggesting less variability in the number of generated triggering inputs across the different runs. These observations are also consistent for the Cifar-10 dataset. These CV values demonstrate the relative stability and reliability of DiffGAN in generating triggering images, with less variability compared to DRfuzz.
@@
Table IV: RQ1 results: Number of generated triggering images for models trained on MNIST dataset using DiffGAN and DRfuzz
@@
Table V: RQ1 results: Number of generated triggering images for models trained on Cifar-10 dataset using DiffGAN and DRfuzz
@@
To systematically compare DiffGAN and DRfuzz in terms of generating valid triggering inputs within a given time budget, we considered the inputs generated in RQ1 by both methods. More specifically, we considered all four model pairs using the MNIST dataset and ran both approaches for two hours on these models. We chose to focus on MNIST in this RQ due to how resource-intensive is the evaluation process. Indeed, validity assessments require significant manual effort, particularly for labeling generated inputs to ensure they are contextually valid. The lower resolution and greater complexity of CIFAR-10 images further complicate this process, as accurately labeling and validating inputs in CIFAR-10 is considerably more challenging compared to MNIST. Given these challenges, we opted to limit our validity study to the MNIST dataset to ensure its feasibility.
As mentioned in RQ1, we ran each method three times to minimize randomness and recorded all images generated by each approach during each run. After completing the runs, for each approach and model pair, we aggregated the images generated from all three runs into a comprehensive dataset. By combining the results from three separate runs for each configuration, we were able to mitigate the effects of randomness in the generation process, thereby increasing the reliability of our comparisons.
Table VI: MNIST Improvement
@@
Table VII: Cifar-10 Improvement
@@
Selecting representative samples. Due to the large number of triggering inputs generated by DiffGAN and DRfuzz, we selected a sample of triggering inputs from each aggregated dataset. For each setting, we determined the sample size that can confidently represent the whole corresponding dataset with a 95% confidence level and a margin of error of 5% [^5] As computed from https://www.calculator.net/sample-size-calculator.html. As a result, in our experiments, the sample size varied from 361 to 381 across the four settings that we considered, as shown in Table VIII. This approach helped choose a minimal subset that was large enough to provide results that accurately reflected the overall distribution of inputs generated by DiffGAN and DRfuzz [62].
Manual validity checking. From the non-redundant set, we randomly sampled images for validation, according to the sample size determined the previous step. To ensure unbiased evaluation, we shuffled images generated by both DiffGAN and DRfuzz and presented them anonymously to three independent expert evaluators. These experts were tasked with assessing the validity of each image based on predefined criteria. Specifically, they labeled images as either valid or invalid. Invalid images were defined as those that displayed two numbers, no number, or other issues such as (1) Images that were a 50/50 blend of two digits and could not be clearly identified as a specific digit, (2) Incomplete or blended digits that could not be classified into one category by consensus, (3) Images containing random shapes or non-digit drawings, or (4) Images that were completely white or dark, with no discernible content. Examples of valid and invalid inputs generated by both approaches are depicted in Figure 4. To determine the validity of each image, we considered the agreement among the evaluators. If at least two out of three evaluators labeled an image as valid, we considered it valid. To quantify the level of agreement between the evaluators, we employ Cohen’s kappa [63]—a statistical measure that accounts for agreement occurring by chance. In our study, the value of Cohen’s kappa is 70%, indicating a high level of agreement among the evaluators and underscoring the reliability of the manual labeling process.
@@
Reporting the total number of valid inputs. We calculated the proportion of valid images for each method by dividing the number of valid images by the total number of sampled images. We then multiplied the validity ratio, calculated from the subset, by the average number of images generated in the runs to estimate the expected number of valid images in each run. Additionally, we determined the 95% confidence interval as reported in Table VIII.
Finally, we compared the expected number of valid inputs generated by DiffGAN and DRfuzz to assess which method was more effective at producing valid triggering inputs within the same time budget of two hours. The results presented in Table VIII, clearly suggest that DiffGAN significantly outperforms DRfuzz in generating more valid triggering inputs. For instance, in the first setting, it is estimated that DiffGAN generated 8,292 valid triggering inputs, whereas DRfuzz produced only 1,999 such inputs. This corresponds to a validity ratio of 91.25% for DiffGAN compared to 82.46% for DRfuzz. Importantly, those results were consistent across all the settings we considered. To further analyze the statistical significance of our results we performed a statistical analysis using the Wilcoxon signed-rank test [65], with a significance level of $\alpha=0.05$, to assess whether DiffGAN significantly outperforms DRfuzz in generating valid triggering inputs. The Wilcoxon signed-rank test, a non-parametric method, is used to compare the medians of continuous variables for paired samples and is ideal for our case since it does not require any assumptions about the data distribution. Instead of considering the number of valid inputs, we relied on the validity ratio due to differences in sample sizes. We evaluated the validity ratios from 4000 samples for DiffGAN and 4000 samples for DRfuzz, selected via bootstrapping across four different scenarios (1000 samples per scenario). The Wilcoxon signed-rank test, with w-statistic = 343465.5 and a p-value of 0.000001, confirmed that DiffGAN significantly outperformed DRfuzz at generating valid triggering inputs across all scenarios.
Table VIII: Validity Comparison
@@
The results, as shown in Tables IX and X, consistently demonstrate that DiffGAN outperforms DRfuzz in generating more diverse triggering inputs, when considering both metrics, highlighting its effectiveness in exploring a broader range of input variations.
Similarly to RQ1, we analysed the variability of our diversity results using the CV score. Our CV analysis shows that both DiffGAN and DRfuzz exhibit low variability across different runs, with DiffGAN demonstrating greater consistency. DRfuzz’s GD diversity has a CV range of 6.76% to 18.44%, while DiffGAN is more stable with a CV range of 2.24% to 14.23%. In terms of Shannon metrics, DRfuzz shows a CV range between 0.09% and 0.70%, whereas DiffGAN is even more consistent, with a CV range of 0.02% to 0.33%. These results underscore the superior stability of DiffGAN’s outputs compared to DRfuzz.
Furthermore, the superior performance of DiffGAN in generating more diverse triggering inputs can be attributed to its large latent space, which allows for a greater variety of generated inputs, and its diversity-based fitness function specifically designed to maximize the diversity of generated inputs, a feature that is limited in DRfuzz. Indeed, DRfuzz focuses on triggering behavioral disagreements through limited mutation rules, which, while effective in some contexts, do not inherently prioritize or ensure diversity. As a result, the input sets generated by DRfuzz tend to be less diverse compared to DiffGAN.
Table IX: Diversity scores of MNIST
@@
Table X: Diversity scores of Cifar-10
@@
We aim in this question to study whether we can better guide the training of an ML-based model selection mechanism that consists of (1) two classifiers (model 1 and model 2) responsible for image classification, and (2) a third model that dynamically selects which of the two classifiers to use based on the input characteristics, which is based on Random Forest (RF). It predicts, based on the characteristics of the incoming input, which of the two classifiers (model 1 or model 2) will produce a more accurate output, without the need to run both models in an online setting.
When properly trained with triggering inputs, the RF model should learn when each classifier is more accurate, effectively serving as a decision mechanism for selecting the optimal classifier for a given input. We choose RF because: (1) it scales effectively to handle a large number of features, and (2) its robustness against overfitting [71, 72, 73]. We do not intend to provide a novel voting mechanism here but, instead, provide an example of an important application scenario for the triggering inputs we generate. It also serves as a basis to compare the value of the inputs generated by DiffGAN to enable such model selection and compare it with that of DRfuzz.
@@
To compare the accuracy of the RF-based model selection systems, we categorized the images into three groups based on whether one model predicted correctly or if both models failed. We excluded images where both models failed from further analysis due to their low occurrence and lack of relevance to the comparison (i.e., if both models fail, it does not matter which classification model is selected). We trained the RF classifiers using features extracted from a pre-trained VGG16 model, rather than raw pixel data, in an attempt to improve accuracy. The VGG16 model provides feature vectors that capture more abstract and informative patterns than pixel-based features, which is expected to improve the RF model’s ability to distinguish between correct and incorrect predictions. For each method and model pair, we selected 700 valid labeled images and used their VGG16 features to train eight RF-based selection models. To ensure a fair comparison between the RF models trained on different datasets in each scenario, we need a common test dataset. For each scenario, we selected 100 images from the DiffGAN-generated test set and 100 from the DRfuzz-generated test set, creating a combined test set of 200 images. This test set allowed us to evaluate the performance of the RF models using the same test set in an unbiased manner. Additionally, we applied the same feature extraction process as for the training sets.
We report the accuracy levels of each of the four selection model we trained in Figure 5. As clearly visible, the RF-based selection models trained using triggering inputs generated with DiffGAN are far more accurate than those trained with inputs from DRfuzz. For instance, in the first scenario involving two RF-based selection models built using classifiers Model 1 and Model 2 from Table I, the RF model trained with DRfuzz achieved an accuracy of 78%, whereas the one trained with DiffGAN reached a much higher accuracy of 87%. Overall, the accuracy levels for the RF models trained with DiffGAN range between 79% and 87% compared to a range of 70% to 78% for those with DRfuzz inputs. This highlights that DiffGAN produces more effective triggering inputs for training, allowing the RF model to better capture the conditions under which one classifier outperforms the other. As a result, RF models trained using DiffGAN inputs are more accurate in selecting the best-performing classifier, leading to a better model selection mechanism.
Figure 5: Accuracy of Random Forest in Selecting the Best Model